\begin{thebibliography}{999}

\bibitem[Aghanim et~al.(2020)]{Planck:2018vyg}
Aghanim, N. et~al. {[Planck Collaboration]} \newblock {Planck 2018 results. VI. Cosmological parameters}.
\newblock {\em Astron. Astrophys.} {\bf 2020}, {\em 641},~A6;
%  \href{http://arxiv.org/abs/1807.06209}{{\normalfont
%  [arXiv:astro-ph.CO/1807.06209]}}.
\newblock Erratum in \emph{Astron. Astrophys.} \textbf{2021}, \emph{652}, C4.
 [\href{http://doi.org/10.1051/0004-6361/201833910}{CrossRef}]

\bibitem[Qu et~al.(2024)]{ACT:2023dou}
Qu, F.J.; Sherwin, B.D.; Madhavacheril, M.S.; Han, D.; Crowley, K.T.; Abril-Cabezas, I.; Ade, P.A.R.; Aiola, S.; Alford, T.; Amiri, M.; et~al. \newblock {The Atacama Cosmology Telescope: A Measurement of the DR6 CMB
Lensing Power Spectrum and Its Implications for Structure Growth}.
\newblock {\em Astrophys. J.} {\bf 2024}, {\em 962},~112.
%  \href{http://arxiv.org/abs/2304.05202}{{\normalfont
%  [arXiv:astro-ph.CO/2304.05202]}}.
[\href{http://dx.doi.org/10.3847/1538-4357/acfe06}{CrossRef}]

\bibitem[Balkenhol et~al.(2023)]{SPT-3G:2022hvq}
Balkenhol, L. et~al. {[SPT-3G Collaboration]} \newblock {Measurement of the CMB temperature power spectrum and constraints on
cosmology from the SPT-3G 2018 TT, TE, and EE dataset}.
\newblock {\em Phys. Rev. D} {\bf 2023}, {\em 108},~023510.
%  \href{http://arxiv.org/abs/2212.05642}{{\normalfont
%  [arXiv:astro-ph.CO/2212.05642]}}.
[\href{http://dx.doi.org/10.1103/PhysRevD.108.023510}{CrossRef}]

\bibitem[Alam et~al.(2021)]{eBOSS:2020yzd}
Alam, S.; Aubert, M.; Avila, S.; Balland, C.; Bautista, J.E.; Bershady, M.A.; Bizyaev, D.; Blanton, M.R.; Bolton, A.S.; \mbox{Bovy, J.; et al.}\newblock {Completed SDSS-IV extended Baryon Oscillation Spectroscopic Survey:
Cosmological implications from two decades of spectroscopic surveys at the
Apache Point Observatory}.
\newblock {\em Phys. Rev. D} {\bf 2021}, {\em 103},~083533.
%  \href{http://arxiv.org/abs/2007.08991}{{\normalfont
%  [arXiv:astro-ph.CO/2007.08991]}}.
[\href{http://dx.doi.org/10.1103/PhysRevD.103.083533}{CrossRef}]

\bibitem[Asgari et~al.(2021)]{KiDS:2020suj}
Asgari, M. et~al. {[KiDS Collaboration]} \newblock {KiDS-1000 Cosmology: Cosmic shear constraints and comparison between
two point statistics}.
\newblock {\em Astron. Astrophys.} {\bf 2021}, {\em 645},~A104.
%  \href{http://arxiv.org/abs/2007.15633}{{\normalfont
%  [arXiv:astro-ph.CO/2007.15633]}}.
[\href{http://dx.doi.org/10.1051/0004-6361/202039070}{CrossRef}]

\bibitem[Abbott et~al.(2022)]{DES:2021wwk}
Abbott, T.M.C. et~al. {[DES Collaboration]} \newblock {Dark Energy Survey Year 3 results: Cosmological constraints from
galaxy clustering and weak lensing}.
\newblock {\em Phys. Rev. D} {\bf 2022}, {\em 105},~023520.
%  \href{http://arxiv.org/abs/2105.13549}{{\normalfont
%  [arXiv:astro-ph.CO/2105.13549]}}.
[\href{http://dx.doi.org/10.1103/PhysRevD.105.023520}{CrossRef}]

\bibitem[Perivolaropoulos and Skara(2022)]{Perivolaropoulos:2021jda}
Perivolaropoulos, L.; Skara, F.
\newblock {Challenges for \ensuremath{\Lambda}CDM: An update}.
\newblock {\em New Astron. Rev.} {\bf 2022}, {\em 95},~101659.
%  \href{http://arxiv.org/abs/2105.05208}{{\normalfont
%  [arXiv:astro-ph.CO/2105.05208]}}.
[\href{http://dx.doi.org/10.1016/j.newar.2022.101659}{CrossRef}]

\bibitem[Abdalla et~al.(2022)]{Abdalla:2022yfr}
Abdalla, E.; Abellán, G.F.; Aboubrahim, A.; Agnello, A.; Akarsu, Ö.; Akrami, Y.; Alestas, G.;  Aloni, D.; Amendola, L.; Anchordoqui, L.A.; et~al. \newblock {Cosmology intertwined: A review of the particle physics,
astrophysics, and cosmology associated with the cosmological tensions and
anomalies}.
\newblock {\em J. High Energy Astrophys.} {\bf 2022}, {\em 34},~49--211.
%  \href{http://arxiv.org/abs/2203.06142}{{\normalfont
%  [arXiv:astro-ph.CO/2203.06142]}}.
[\href{http://dx.doi.org/10.1016/j.jheap.2022.04.002}{CrossRef}]

\bibitem[Verde et~al.(2019)Verde, Treu, and Riess]{Verde:2019ivm}
Verde, L.; Treu, T.; Riess, A.G.
\newblock {Tensions between the Early and the Late Universe}.
\newblock {\em Nat. Astron.} {\bf 2019}, {\em 3},~891.
%  \href{http://arxiv.org/abs/1907.10625}{{\normalfont
%  [arXiv:astro-ph.CO/1907.10625]}}.
[\href{http://dx.doi.org/10.1038/s41550-019-0902-0}{CrossRef}]

\bibitem[Di~Valentino et~al.(2021{\natexlab{a}})]{DiValentino:2020zio}
Di Valentino, E.; Anchordoqui, L.A.; Akarsu, Ö.; Ali-Haimoud, Y.; Amendola, L.; Arendse, N.; Asgari, M.; Ballardini, M.; Basilakos, S.; Battistelli, E.; et~al.
\newblock {Snowmass2021---Letter of interest cosmology intertwined II: The
hubble constant tension}.
\newblock {\em Astropart. Phys.} {\bf 2021}, {\em 131},~102605.
%  \href{http://arxiv.org/abs/2008.11284}{{\normalfont
%  [arXiv:astro-ph.CO/2008.11284]}}.
[\href{http://dx.doi.org/10.1016/j.astropartphys.2021.102605}{CrossRef}]

\bibitem[Di~Valentino et~al.(2021{\natexlab{b}})Di~Valentino, Mena, Pan,
Visinelli, Yang, Melchiorri, Mota, Riess, and Silk]{DiValentino:2021izs}
Di~Valentino, E.; Mena, O.; Pan, S.; Visinelli, L.; Yang, W.; Melchiorri, A.;
Mota, D.F.; Riess, A.G.; Silk, J.
\newblock {In the realm of the Hubble tension---A review of
solutions}.
\newblock {\em Class. Quant. Grav.} {\bf 2021}, {\em 38},~153001.
%  \href{http://arxiv.org/abs/2103.01183}{{\normalfont
%  [arXiv:astro-ph.CO/2103.01183]}}.
[\href{http://dx.doi.org/10.1088/1361-6382/ac086d}{CrossRef}]

\bibitem[Vagnozzi(2023)]{Vagnozzi:2023nrq}
Vagnozzi, S.
\newblock {Seven Hints That Early-Time New Physics Alone Is Not Sufficient to
Solve the Hubble Tension}.
\newblock {\em Universe} {\bf 2023}, {\em 9},~393.
%  \href{http://arxiv.org/abs/2308.16628}{{\normalfont
%  [arXiv:astro-ph.CO/2308.16628]}}.
[\href{http://dx.doi.org/10.3390/universe9090393}{CrossRef}]

\bibitem[Dainotti et~al.(2023)Dainotti, De~Simone, Montani;, Schiavone;, and
Lambiase.]{Dainotti:2023yrk}
Dainotti, M.; De~Simone, B.; Montani, G.; Schiavone, T.; Lambiase, G.
\newblock {The Hubble constant tension: Current status and future perspectives
through new cosmological probes}.
\newblock {\em PoS} {\bf 2023}, {\em CORFU2022},~235.
%  \href{http://arxiv.org/abs/2301.10572}{{\normalfont
%  [arXiv:astro-ph.CO/2301.10572]}}.
[\href{http://dx.doi.org/10.22323/1.436.0235}{CrossRef}]

\bibitem[Akarsu et~al.(2024)Akarsu, Colg\'ain, Sen, and
Sheikh-Jabbari]{Akarsu:2024qiq}
Akarsu, O.; Colg\'ain, E.O.; Sen, A.A.; Sheikh-Jabbari, M.M.
\newblock {$\Lambda$CDM Tensions: Localising Missing Physics through
Consistency Checks}. {\emph{arXiv} \textbf{2024}, arXiv:2402.04767.} 

% We confirm and agree.

%\newblock  \href{http://arxiv.org/abs/2402.04767}{{\normalfont
%  [arXiv:astro-ph.CO/2402.04767]}}.

\bibitem[Kamionkowski and Riess(2023)]{Kamionkowski:2022pkx}
Kamionkowski, M.; Riess, A.G.
\newblock {The Hubble Tension and Early Dark Energy}.
\newblock {\em Ann. Rev. Nucl. Part. Sci.} {\bf 2023}, {\em 73},~153--180.
%  \href{http://arxiv.org/abs/2211.04492}{{\normalfont
%  [arXiv:astro-ph.CO/2211.04492]}}.
[\href{http://dx.doi.org/10.1146/annurev-nucl-111422-024107}{CrossRef}]

\bibitem[Dainotti et~al.(2021)Dainotti, De~Simone, Schiavone, Montani, Rinaldi,
and Lambiase]{Dainotti:2021pqg}
Dainotti, M.G.; De~Simone, B.; Schiavone, T.; Montani, G.; Rinaldi, E.;
Lambiase, G.
\newblock {On the Hubble constant tension in the SNe Ia Pantheon sample}.
\newblock {\em Astrophys. J.} {\bf 2021}, {\em 912},~150.
%  \href{http://arxiv.org/abs/2103.02117}{{\normalfont
%  [arXiv:astro-ph.CO/2103.02117]}}.
[\href{http://dx.doi.org/10.3847/1538-4357/abeb73}{CrossRef}]

\bibitem[Jedamzik et~al.(2021)Jedamzik, Pogosian, and Zhao]{Jedamzik:2020zmd}
Jedamzik, K.; Pogosian, L.; Zhao, G.B.
\newblock {Why reducing the cosmic sound horizon alone can not fully resolve
the Hubble tension}.
\newblock {\em Commun. Phys.} {\bf 2021}, {\em 4},~123.
%  \href{http://arxiv.org/abs/2010.04158}{{\normalfont
%  [arXiv:astro-ph.CO/2010.04158]}}.
[\href{http://dx.doi.org/10.1038/s42005-021-00628-x}{CrossRef}]

\bibitem[Vagnozzi et~al.(2022)Vagnozzi, Pacucci, and Loeb]{Vagnozzi:2021tjv}
Vagnozzi, S.; Pacucci, F.; Loeb, A.
\newblock {Implications for the Hubble tension from the ages of the oldest
astrophysical objects}.
\newblock {\em J. High Energy Astrophys.} {\bf 2022}, {\em 36},~27--35.
%  \href{http://arxiv.org/abs/2105.10421}{{\normalfont
%  [arXiv:astro-ph.CO/2105.10421]}}.
[\href{http://dx.doi.org/10.1016/j.jheap.2022.07.004}{CrossRef}]

\bibitem[Di~Valentino et~al.(2021)]{DiValentino:2020vvd}
Di Valentino, E.; Anchordoqui, L.A.; Akarsu, Ö.; Ali-Haimoud, Y.; Amendola, L.; Arendse, N.; Asgari, M.; Ballardini, M.; Basilakos, S.; Battistelli, E.; et~al.
\newblock {Cosmology Intertwined III: $f \sigma_8$ and $S_8$}.
\newblock {\em Astropart. Phys.} {\bf 2021}, {\em 131},~102604.
%  \href{http://arxiv.org/abs/2008.11285}{{\normalfont
%  [arXiv:astro-ph.CO/2008.11285]}}.
[\href{http://dx.doi.org/10.1016/j.astropartphys.2021.102604}{CrossRef}]

\bibitem[Philcox and Ivanov(2022)]{Philcox:2021kcw}
Philcox, O.H.E.; Ivanov, M.M.
\newblock {BOSS DR12 full-shape cosmology: \ensuremath{\Lambda}CDM constraints
from the large-scale galaxy power spectrum and bispectrum monopole}.
\newblock {\em Phys. Rev. D} {\bf 2022}, {\em 105},~043517.
%  \href{http://arxiv.org/abs/2112.04515}{{\normalfont
%  [arXiv:astro-ph.CO/2112.04515]}}.
[\href{http://dx.doi.org/10.1103/PhysRevD.105.043517}{CrossRef}]

\bibitem[Heisenberg et~al.(2022)Heisenberg, Villarrubia-Rojo, and
Zosso]{Heisenberg:2022gqk}
Heisenberg, L.; Villarrubia-Rojo, H.; Zosso, J.
\newblock {Can late-time extensions solve the $H_0$ and $\sigma_8$
tensions?}
\newblock {\em Phys. Rev. D} {\bf 2022}, {\em 106},~043503.
%  \href{http://arxiv.org/abs/2202.01202}{{\normalfont
%  [arXiv:astro-ph.CO/2202.01202]}}.
[\href{http://dx.doi.org/10.1103/PhysRevD.106.043503}{CrossRef}]

\bibitem[Benisty(2021)]{Benisty:2020kdt}
Benisty, D.
\newblock {Quantifying the $S_8$ tension with the Redshift Space Distortion
data set}.
\newblock {\em Phys. Dark Univ.} {\bf 2021}, {\em 31},~100766.
%  \href{http://arxiv.org/abs/2005.03751}{{\normalfont
%  [arXiv:astro-ph.CO/2005.03751]}}.
[\href{http://dx.doi.org/10.1016/j.dark.2020.100766}{CrossRef}]

\bibitem[Abbott et~al.(2023)]{Kilo-DegreeSurvey:2023gfr}
Abbott, T.M.C. et~al. {[Dark Energy Survey and Kilo-Degree Survey Collaboration]} \newblock {DES Y3 + KiDS-1000: Consistent cosmology combining cosmic shear
surveys}.
\newblock {\em Open J. Astrophys.} {\bf 2023}, {\em 6},~1--40.
%  \href{http://arxiv.org/abs/2305.17173}{{\normalfont
%  [arXiv:astro-ph.CO/2305.17173]}}.
[\href{http://dx.doi.org/10.21105/astro.2305.17173}{CrossRef}]

\bibitem[Tr\"oster et~al.(2021)]{KiDS:2020ghu}
Tr\"oster, T. et~al. {[KiDS Collaboration]} \newblock {KiDS-1000 Cosmology: Constraints beyond flat
\ensuremath{\Lambda}CDM}.
\newblock {\em Astron. Astrophys.} {\bf 2021}, {\em 649},~A88.
%  \href{http://arxiv.org/abs/2010.16416}{{\normalfont
%  [arXiv:astro-ph.CO/2010.16416]}}.
[\href{http://dx.doi.org/10.1051/0004-6361/202039805}{CrossRef}]

\bibitem[Joudaki et~al.(2020)]{Joudaki:2019pmv}
Joudaki, S.; Hildebrandt, H.; Traykova, D.; Chisari, N.E.; Heymans, C.; Kannawadi, A.; Kuijken, K.; Wright, A.H.; Asgari, M.; Erben, T.; et~al.
\newblock {KiDS+VIKING-450 and DES-Y1 combined: Cosmology with cosmic shear}.
\newblock {\em Astron. Astrophys.} {\bf 2020}, {\em 638},~L1.
%  \href{http://arxiv.org/abs/1906.09262}{{\normalfont
%  [arXiv:astro-ph.CO/1906.09262]}}.
[\href{http://dx.doi.org/10.1051/0004-6361/201936154}{CrossRef}]

\bibitem[Tr\"oster et~al.(2020)]{Troster:2019ean}
Tr\"oster, T.; Sánchez, A.G.; Asgari, M.; Blake, C.; Crocce, M.; Heymans, C.; Hildebrandt,  H.; Joachimi, B.; Joudaki, S.; Kannawadi, A.; et~al. \newblock {Cosmology from large-scale structure: Constraining $\Lambda$CDM with
BOSS}.
\newblock {\em Astron. Astrophys.} {\bf 2020}, {\em 633},~L10.
%  \href{http://arxiv.org/abs/1909.11006}{{\normalfont
%  [arXiv:astro-ph.CO/1909.11006]}}.
[\href{http://dx.doi.org/10.1051/0004-6361/201936772}{CrossRef}]

\bibitem[Aluri et~al.(2023)]{Aluri:2022hzs}
Aluri, P.K.; Cea, P.; Chingangbam, P.; Chu, M.C.; Clowes, R.G.; Hutsemékers, D.; Kochappan, J.P.; Lopez, A.M.; Liu, L.; Martens, N.C.; et~al. \newblock {Is the observable Universe consistent with the cosmological
principle?}
\newblock {\em Class. Quant. Grav.} {\bf 2023}, {\em 40},~094001.
%  \href{http://arxiv.org/abs/2207.05765}{{\normalfont
%  [arXiv:astro-ph.CO/2207.05765]}}.
[\href{http://dx.doi.org/10.1088/1361-6382/acbefc}{CrossRef}]

\bibitem[Secrest et~al.(2021)Secrest, von Hausegger, Rameez, Mohayaee, Sarkar,
and Colin]{Secrest:2020has}
Secrest, N.J.; von Hausegger, S.; Rameez, M.; Mohayaee, R.; Sarkar, S.; Colin,
J.
\newblock {A Test of the Cosmological Principle with Quasars}.
\newblock {\em Astrophys. J. Lett.} {\bf 2021}, {\em 908},~L51.
%  \href{http://arxiv.org/abs/2009.14826}{{\normalfont
%  [arXiv:astro-ph.CO/2009.14826]}}.
[\href{http://dx.doi.org/10.3847/2041-8213/abdd40}{CrossRef}]

\bibitem[Uzan et~al.(2008)Uzan, Clarkson, and Ellis]{Uzan:2008qp}
Uzan, J.P.; Clarkson, C.; Ellis, G.F.R.
\newblock {Time drift of cosmological redshifts as a test of the Copernican
principle}.
\newblock {\em Phys. Rev. Lett.} {\bf 2008}, {\em 100},~191303.
%  \href{http://arxiv.org/abs/0801.0068}{{\normalfont
%  [arXiv:astro-ph/0801.0068]}}.
[\href{http://dx.doi.org/10.1103/PhysRevLett.100.191303}{CrossRef}]

\bibitem[Mc~Conville and Colg\'ain(2023)]{McConville:2023xav}
Mc~Conville, R.; Colg\'ain, E.O.
\newblock {Anisotropic distance ladder in Pantheon+supernovae}.
\newblock {\em Phys. Rev. D} {\bf 2023}, {\em 108},~123533.
%  \href{http://arxiv.org/abs/2304.02718}{{\normalfont
%  [arXiv:astro-ph.CO/2304.02718]}}.
[\href{http://dx.doi.org/10.1103/PhysRevD.108.123533}{CrossRef}]

\bibitem[Watkins et~al.(2023)Watkins, Allen, Bradford, Ramon, Walker, Feldman,
Cionitti, Al-Shorman, Kourkchi, and Tully]{Watkins:2023rll}
Watkins, R.; Allen, T.; Bradford, C.J.; Ramon, A.; Walker, A.; Feldman, H.A.;
Cionitti, R.; Al-Shorman, Y.; Kourkchi, E.; Tully, R.B.
\newblock {Analysing the large-scale bulk flow using cosmicflows4: Increasing
tension with the standard cosmological model}.
\newblock {\em Mon. Not. R. Astron. Soc.} {\bf 2023}, {\em 524},~1885--1892.
%  \href{http://arxiv.org/abs/2302.02028}{{\normalfont
%  [arXiv:astro-ph.CO/2302.02028]}}.
[\href{http://dx.doi.org/10.1093/mnras/stad1984}{CrossRef}]

\bibitem[Cowell et~al.(2023)Cowell, Dhawan, and Macpherson]{Cowell:2022ehf}
Cowell, J.A.; Dhawan, S.; Macpherson, H.J.
\newblock {Potential signature of a quadrupolar hubble expansion in
Pantheon+supernovae}.
\newblock {\em Mon. Not. R. Astron. Soc.} {\bf 2023}, {\em 526},~1482--1494.
%  \href{http://arxiv.org/abs/2212.13569}{{\normalfont
%  [arXiv:astro-ph.CO/2212.13569]}}.
[\href{http://dx.doi.org/10.1093/mnras/stad2788}{CrossRef}]

\bibitem[Dam et~al.(2023)Dam, Lewis, and Brewer]{Dam:2022wwh}
Dam, L.; Lewis, G.F.; Brewer, B.J.
\newblock {Testing the cosmological principle with CatWISE quasars: A bayesian
analysis of the number-count dipole}.
\newblock {\em Mon. Not. R. Astron. Soc.} {\bf 2023}, {\em 525},~231--245.
%  \href{http://arxiv.org/abs/2212.07733}{{\normalfont
%  [arXiv:astro-ph.CO/2212.07733]}}.
[\href{http://dx.doi.org/10.1093/mnras/stad2322}{CrossRef}]

\bibitem[Krishnan et~al.(2023)Krishnan, Mondol, and
Sheikh-Jabbari]{Krishnan:2022qbv}
Krishnan, C.; Mondol, R.; Sheikh-Jabbari, M.M.
\newblock {Dipole cosmology: The Copernican paradigm beyond FLRW}.
\newblock {\em J. Cosmol. Astropart. Phys.} {\bf 2023}, {\em 07},~020.
%  \href{http://arxiv.org/abs/2209.14918}{{\normalfont
%  [arXiv:astro-ph.CO/2209.14918]}}.
[\href{http://dx.doi.org/10.1088/1475-7516/2023/07/020}{CrossRef}]

\bibitem[Bengaly et~al.(2024)Bengaly, Pigozzo, and Alcaniz]{Bengaly:2024ree}
Bengaly, C.A.P.; Pigozzo, C.; Alcaniz, J.S.
\newblock {Testing the isotropy of cosmic acceleration with Pantheon+SH0ES: A
cosmographic analysis}. {\emph{arXiv} {\bf 2024}, arXiv:2402.1774.}
%\newblock  \href{http://arxiv.org/abs/2402.17741}{{\normalfont
%  [arXiv:astro-ph.CO/2402.17741]}}.

\bibitem[Ebrahimian et~al.(2023)Ebrahimian, Krishnan, Mondol, and
Sheikh-Jabbari]{Ebrahimian:2023svi}
Ebrahimian, E.; Krishnan, C.; Mondol, R.; Sheikh-Jabbari, M.M.
\newblock {Towards A Realistic Dipole Cosmology: The Dipole $\Lambda$CDM Model}.
{\emph{arXiv} {\bf 2023}, arXiv:2305.16177.}
%\newblock  \href{http://arxiv.org/abs/2305.16177}{{\normalfont
%  [arXiv:astro-ph.CO/2305.16177]}}.

\bibitem[Perivolaropoulos(2023)]{Perivolaropoulos:2023tdt}
Perivolaropoulos, L.
\newblock {Isotropy properties of the absolute luminosity magnitudes of SnIa in
the Pantheon+ and SH0ES samples}.
\newblock {\em Phys. Rev. D} {\bf 2023}, {\em 108},~063509.
%  \href{http://arxiv.org/abs/2305.12819}{{\normalfont
%  [arXiv:astro-ph.CO/2305.12819]}}.
[\href{http://dx.doi.org/10.1103/PhysRevD.108.063509}{CrossRef}]

\bibitem[Bahamonde et~al.(2023)Bahamonde, Dialektopoulos, Escamilla-Rivera,
Farrugia, Gakis, Hendry, Hohmann, Levi~Said, Mifsud, and
Di~Valentino]{Bahamonde:2021gfp}
Bahamonde, S.; Dialektopoulos, K.F.; Escamilla-Rivera, C.; Farrugia, G.; Gakis,
V.; Hendry, M.; Hohmann, M.; Levi~Said, J.; Mifsud, J.; Di~Valentino, E.
\newblock {Teleparallel gravity: From theory to cosmology}.
\newblock {\em Rept. Prog. Phys.} {\bf 2023}, {\em 86},~026901.
%  \href{http://arxiv.org/abs/2106.13793}{{\normalfont
%  [arXiv:gr-qc/2106.13793]}}.
[\href{http://dx.doi.org/10.1088/1361-6633/ac9cef}{CrossRef}] [\href{http://www.ncbi.nlm.nih.gov/pubmed/36279849}{PubMed}]

\bibitem[Adi and Kovetz(2021)]{Adi:2020qqf}
Adi, T.; Kovetz, E.D.
\newblock {Can conformally coupled modified gravity solve the Hubble tension?}
\newblock {\em Phys. Rev. D} {\bf 2021}, {\em 103},~023530.
%  \href{http://arxiv.org/abs/2011.13853}{{\normalfont
%  [arXiv:astro-ph.CO/2011.13853]}}.
[\href{http://dx.doi.org/10.1103/PhysRevD.103.023530}{CrossRef}]

\bibitem[Marra and Perivolaropoulos(2021)]{Marra:2021fvf}
Marra, V.; Perivolaropoulos, L.
\newblock {Rapid transition of Geff at zt\ensuremath{\simeq}0.01 as a possible
solution of the Hubble and growth tensions}.
\newblock {\em Phys. Rev. D} {\bf 2021}, {\em 104},~L021303.
%  \href{http://arxiv.org/abs/2102.06012}{{\normalfont
%  [arXiv:astro-ph.CO/2102.06012]}}.
[\href{http://dx.doi.org/10.1103/PhysRevD.104.L021303}{CrossRef}]

\bibitem[Odintsov et~al.(2021)Odintsov, S\'aez-Chill\'on~G\'omez, and
Sharov]{Odintsov:2020qzd}
Odintsov, S.D.; S\'aez-Chill\'on~G\'omez, D.; Sharov, G.S.
\newblock {Analyzing the $H_0$ tension in $F(R)$ gravity models}.
\newblock {\em Nucl. Phys. B} {\bf 2021}, {\em 966},~115377.
%  \href{http://arxiv.org/abs/2011.03957}{{\normalfont
%  [arXiv:gr-qc/2011.03957]}}.
[\href{http://dx.doi.org/10.1016/j.nuclphysb.2021.115377}{CrossRef}]

\bibitem[Ballardini et~al.(2020)Ballardini, Braglia, Finelli, Paoletti,
Starobinsky, and Umilt\`a]{Ballardini:2020iws}
Ballardini, M.; Braglia, M.; Finelli, F.; Paoletti, D.; Starobinsky, A.A.;
Umilt\`a, C.
\newblock {Scalar-tensor theories of gravity, neutrino physics, and the $H_0$
tension}.
\newblock {\em J. Cosmol. Astropart. Phys.} {\bf 2020}, {\em 10},~044.
%  \href{http://arxiv.org/abs/2004.14349}{{\normalfont
%  [arXiv:astro-ph.CO/2004.14349]}}.
[\href{http://dx.doi.org/10.1088/1475-7516/2020/10/044}{CrossRef}]

\bibitem[Wang and Mota(2020)]{Wang:2020zfv}
Wang, D.; Mota, D.
\newblock {Can $f(T)$ gravity resolve the $H_0$ tension?}
\newblock {\em Phys. Rev. D} {\bf 2020}, {\em 102},~063530.
%  \href{http://arxiv.org/abs/2003.10095}{{\normalfont
%  [arXiv:astro-ph.CO/2003.10095]}}.
[\href{http://dx.doi.org/10.1103/PhysRevD.102.063530}{CrossRef}]

\bibitem[Skara and Perivolaropoulos(2020)]{Skara:2019usd}
Skara, F.; Perivolaropoulos, L.
\newblock {Tension of the $E_G$ statistic and redshift space distortion data
with the Planck---$\Lambda CDM$ model and implications for weakening
gravity}.
\newblock {\em Phys. Rev. D} {\bf 2020}, {\em 101},~063521.
%  \href{http://arxiv.org/abs/1911.10609}{{\normalfont
%  [arXiv:astro-ph.CO/1911.10609]}}.
[\href{http://dx.doi.org/10.1103/PhysRevD.101.063521}{CrossRef}]

\bibitem[Escamilla-Rivera and Levi~Said(2020)]{Escamilla-Rivera:2019ulu}
Escamilla-Rivera, C.; Levi~Said, J.
\newblock {Cosmological viable models in $f(T,B)$ theory as solutions to the
$H_0$ tension}.
\newblock {\em Class. Quant. Grav.} {\bf 2020}, {\em 37},~165002.
%  \href{http://arxiv.org/abs/1909.10328}{{\normalfont
%  [arXiv:gr-qc/1909.10328]}}.
[\href{http://dx.doi.org/10.1088/1361-6382/ab939c}{CrossRef}]

\bibitem[Cai et~al.(2020)Cai, Khurshudyan, and Saridakis]{Cai:2019bdh}
Cai, Y.F.; Khurshudyan, M.; Saridakis, E.N.
\newblock {Model-independent reconstruction of $f(T)$ gravity from Gaussian
Processes}.
\newblock {\em Astrophys. J.} {\bf 2020}, {\em 888},~62.
%  \href{http://arxiv.org/abs/1907.10813}{{\normalfont
%  [arXiv:astro-ph.CO/1907.10813]}}.
[\href{http://dx.doi.org/10.3847/1538-4357/ab5a7f}{CrossRef}]

\bibitem[Kazantzidis and Perivolaropoulos(2019)]{Kazantzidis:2019nuh}
Kazantzidis, L.; Perivolaropoulos, L.
\newblock {\({\sigma }_{8}\) Tension. Is Gravity Getting Weaker at Low z?
Observational Evidence and Theoretical Implications}. {\emph{arXiv} {\bf 2019}, arXiv:1907.03176.}
%\newblock  \href{http://arxiv.org/abs/1907.03176}{{\normalfont
%  [arXiv:astro-ph.CO/1907.03176]}}.
%\newblock {\url{https://doi.org/10.1007/978-3-030-83715-0_33}}.
[\href{https://doi.org/10.1007/978-3-030-83715-0_33}{CrossRef}]

\bibitem[Kazantzidis and Perivolaropoulos(2018)]{Kazantzidis:2018rnb}
Kazantzidis, L.; Perivolaropoulos, L.
\newblock {Evolution of the $f\sigma_8$ tension with the Planck15/$\Lambda$CDM
determination and implications for modified gravity theories}.
\newblock {\em Phys. Rev. D} {\bf 2018}, {\em 97},~103503.
%  \href{http://arxiv.org/abs/1803.01337}{{\normalfont
%  [arXiv:astro-ph.CO/1803.01337]}}.
[\href{http://dx.doi.org/10.1103/PhysRevD.97.103503}{CrossRef}]

\bibitem[Nunes(2018)]{Nunes:2018xbm}
Nunes, R.C.
\newblock {Structure formation in $f(T)$ gravity and a solution for $H_0$
tension}.
\newblock {\em J. Cosmol. Astropart. Phys.} {\bf 2018}, {\em 05},~052.
%  \href{http://arxiv.org/abs/1802.02281}{{\normalfont
%  [arXiv:gr-qc/1802.02281]}}.
[\href{http://dx.doi.org/10.1088/1475-7516/2018/05/052}{CrossRef}]

\bibitem[Gangopadhyay et~al.(2023)Gangopadhyay, Pacif, Sami, and
Sharma]{Gangopadhyay:2022bsh}
Gangopadhyay, M.R.; Pacif, S.K.J.; Sami, M.; Sharma, M.K.
\newblock {Generic Modification of Gravity, Late Time Acceleration and Hubble
Tension}.
\newblock {\em Universe} {\bf 2023}, {\em 9},~83.
%  \href{http://arxiv.org/abs/2211.12041}{{\normalfont
%  [arXiv:gr-qc/2211.12041]}}.
[\href{http://dx.doi.org/10.3390/universe9020083}{CrossRef}]

\bibitem[Zumalacarregui(2020)]{Zumalacarregui:2020cjh}
Zumalacarregui, M.
\newblock {Gravity in the Era of Equality: Towards solutions to the Hubble
problem without fine-tuned initial conditions}.
\newblock {\em Phys. Rev. D} {\bf 2020}, {\em 102},~023523.
%  \href{http://arxiv.org/abs/2003.06396}{{\normalfont
%  [arXiv:astro-ph.CO/2003.06396]}}.
[\href{http://dx.doi.org/10.1103/PhysRevD.102.023523}{CrossRef}]

\bibitem[Benevento et~al.(2022)Benevento, Kable, Addison, and
Bennett]{Benevento:2022cql}
Benevento, G.; Kable, J.A.; Addison, G.E.; Bennett, C.L.
\newblock {An Exploration of an Early Gravity Transition in Light of
Cosmological Tensions}.
\newblock {\em Astrophys. J.} {\bf 2022}, {\em 935},~156.
%  \href{http://arxiv.org/abs/2202.09356}{{\normalfont
%  [arXiv:astro-ph.CO/2202.09356]}}.
[\href{http://dx.doi.org/10.3847/1538-4357/ac80fd}{CrossRef}]

\bibitem[Braglia et~al.(2021)Braglia, Ballardini, Finelli, and
Koyama]{Braglia:2020auw}
Braglia, M.; Ballardini, M.; Finelli, F.; Koyama, K.
\newblock {Early modified gravity in light of the $H_0$ tension and LSS data}.
\newblock {\em Phys. Rev. D} {\bf 2021}, {\em 103},~043528.
%  \href{http://arxiv.org/abs/2011.12934}{{\normalfont
%  [arXiv:astro-ph.CO/2011.12934]}}.
[\href{http://dx.doi.org/10.1103/PhysRevD.103.043528}{CrossRef}]

\bibitem[Farhang and Khosravi(2021)]{Farhang:2020sij}
Farhang, M.; Khosravi, N.
\newblock {Phenomenological Gravitational Phase Transition: Reconciliation
between the Late and Early Universe}.
\newblock {\em Phys. Rev. D} {\bf 2021}, {\em 103},~083523.
%  \href{http://arxiv.org/abs/2011.08050}{{\normalfont
%  [arXiv:astro-ph.CO/2011.08050]}}.
[\href{http://dx.doi.org/10.1103/PhysRevD.103.083523}{CrossRef}]

\bibitem[Heisenberg and Villarrubia-Rojo(2021)]{Heisenberg:2020xak}
Heisenberg, L.; Villarrubia-Rojo, H.
\newblock {Proca in the sky}.
\newblock {\em J. Cosmol. Astropart. Phys.} {\bf 2021}, {\em 03},~032.
%  \href{http://arxiv.org/abs/2010.00513}{{\normalfont
%  [arXiv:astro-ph.CO/2010.00513]}}.
[\href{http://dx.doi.org/10.1088/1475-7516/2021/03/032}{CrossRef}]

\bibitem[Ballesteros et~al.(2020)Ballesteros, Notari, and
Rompineve]{Ballesteros:2020sik}
Ballesteros, G.; Notari, A.; Rompineve, F.
\newblock {The $H_0$ tension: $\Delta G_N$ vs. $\Delta N_{\rm eff}$}.
\newblock {\em J. Cosmol. Astropart. Phys.} {\bf 2020}, {\em 11},~024.
%  \href{http://arxiv.org/abs/2004.05049}{{\normalfont
%  [arXiv:astro-ph.CO/2004.05049]}}.
[\href{http://dx.doi.org/10.1088/1475-7516/2020/11/024}{CrossRef}]

\bibitem[Franco~Abell\'an et~al.(2023)Franco~Abell\'an, Braglia, Ballardini,
Finelli, and Poulin]{FrancoAbellan:2023gec}
Franco~Abell\'an, G.; Braglia, M.; Ballardini, M.; Finelli, F.; Poulin, V.
\newblock {Probing early modification of gravity with Planck, ACT and SPT}.
\newblock {\em J. Cosmol. Astropart. Phys.} {\bf 2023}, {\em 12},~017.
%  \href{http://arxiv.org/abs/2308.12345}{{\normalfont
%  [arXiv:astro-ph.CO/2308.12345]}}.
[\href{http://dx.doi.org/10.1088/1475-7516/2023/12/017}{CrossRef}]

\bibitem[Akrami et~al.(2021)]{CANTATA:2021ktz}
Akrami, Y.;  Bahamonde, S.; Luis Blázquez-Salcedo, J.; Böhmer, C.; Bonvin, C.; Bouhmadi-López, M.; Brax, P.; Calcagni, G.; Capozziello, S.; Casa, R.; et~al. \newblock {\em Modified Gravity and Cosmology: An Update by the CANTATA
Network}; {Springer: New York, NY, USA}, 2021. 
%  \href{http://arxiv.org/abs/2105.12582}{{\normalfont
%  [arXiv:gr-qc/2105.12582]}}.
[\href{http://dx.doi.org/10.1007/978-3-030-83715-0}{CrossRef}]

\bibitem[Langlois et~al.(2018)Langlois, Saito, Yamauchi, and
Noui]{Langlois:2017dyl}
Langlois, D.; Saito, R.; Yamauchi, D.; Noui, K.
\newblock {Scalar-tensor theories and modified gravity in the wake of
GW170817}.
\newblock {\em Phys. Rev. D} {\bf 2018}, {\em 97},~061501.
%  \href{http://arxiv.org/abs/1711.07403}{{\normalfont
%  [arXiv:gr-qc/1711.07403]}}.
[\href{http://dx.doi.org/10.1103/PhysRevD.97.061501}{CrossRef}]

\bibitem[Nojiri et~al.(2017)Nojiri, Odintsov, and Oikonomou]{Nojiri:2017ncd}
Nojiri, S.; Odintsov, S.D.; Oikonomou, V.K.
\newblock {Modified Gravity Theories on a Nutshell: Inflation, Bounce and
Late-time Evolution}.
\newblock {\em Phys. Rept.} {\bf 2017}, {\em 692},~1--104.
%  \href{http://arxiv.org/abs/1705.11098}{{\normalfont
%  [arXiv:gr-qc/1705.11098]}}.
[\href{http://dx.doi.org/10.1016/j.physrep.2017.06.001}{CrossRef}]

\bibitem[Joyce et~al.(2016)Joyce, Lombriser, and Schmidt]{Joyce:2016vqv}
Joyce, A.; Lombriser, L.; Schmidt, F.
\newblock {Dark Energy Versus Modified Gravity}.
\newblock {\em Ann. Rev. Nucl. Part. Sci.} {\bf 2016}, {\em 66},~95--122.
%  \href{http://arxiv.org/abs/1601.06133}{{\normalfont
%  [arXiv:astro-ph.CO/1601.06133]}}.
[\href{http://dx.doi.org/10.1146/annurev-nucl-102115-044553}{CrossRef}]

\bibitem[Joudaki et~al.(2022)Joudaki, Ferreira, Lima, and
Winther]{Joudaki:2020shz}
Joudaki, S.; Ferreira, P.G.; Lima, N.A.; Winther, H.A.
\newblock {Testing gravity on cosmic scales: A case study of Jordan-Brans-Dicke
theory}.
\newblock {\em Phys. Rev. D} {\bf 2022}, {\em 105},~043522.
%  \href{http://arxiv.org/abs/2010.15278}{{\normalfont
%  [arXiv:astro-ph.CO/2010.15278]}}.
[\href{http://dx.doi.org/10.1103/PhysRevD.105.043522}{CrossRef}]

\bibitem[Koyama(2016)]{Koyama:2015vza}
Koyama, K.
\newblock {Cosmological Tests of Modified Gravity}.
\newblock {\em Rept. Prog. Phys.} {\bf 2016}, {\em 79},~046902.
%  \href{http://arxiv.org/abs/1504.04623}{{\normalfont
%  [arXiv:astro-ph.CO/1504.04623]}}.
[\href{http://dx.doi.org/10.1088/0034-4885/79/4/046902}{CrossRef}] [\href{http://www.ncbi.nlm.nih.gov/pubmed/27007681}{PubMed}]

\bibitem[Kobayashi(2019)]{Kobayashi:2019hrl}
Kobayashi, T.
\newblock {Horndeski theory and beyond: A review}.
\newblock {\em Rept. Prog. Phys.} {\bf 2019}, {\em 82},~086901.
%  \href{http://arxiv.org/abs/1901.07183}{{\normalfont
%  [arXiv:gr-qc/1901.07183]}}.
[\href{http://dx.doi.org/10.1088/1361-6633/ab2429}{CrossRef}] [\href{http://www.ncbi.nlm.nih.gov/pubmed/31121569}{PubMed}]

\bibitem[Ishak(2019)]{Ishak:2018his}
Ishak, M.
\newblock {Testing General Relativity in Cosmology}.
\newblock {\em Living Rev. Rel.} {\bf 2019}, {\em 22},~1.
%  \href{http://arxiv.org/abs/1806.10122}{{\normalfont
%  [arXiv:astro-ph.CO/1806.10122]}}.
[\href{http://dx.doi.org/10.1007/s41114-018-0017-4}{CrossRef}]

\bibitem[Sakstein and Jain(2017)]{Sakstein:2017xjx}
Sakstein, J.; Jain, B.
\newblock {Implications of the Neutron Star Merger GW170817 for Cosmological
Scalar-Tensor Theories}.
\newblock {\em Phys. Rev. Lett.} {\bf 2017}, {\em 119},~251303.
%  \href{http://arxiv.org/abs/1710.05893}{{\normalfont
%  [arXiv:astro-ph.CO/1710.05893]}}.
[\href{http://dx.doi.org/10.1103/PhysRevLett.119.251303}{CrossRef}] [\href{http://www.ncbi.nlm.nih.gov/pubmed/29303345}{PubMed}]

\bibitem[Amendola et~al.(2018)]{Amendola:2016saw}
Amendola, L.; Appleby, S.; Avgoustidis, A.;  Bacon, D.; Baker, T.; Baldi, M.; Bartolo, N.; Blanchard, A.; Bonvin, C.; Borgani, S.; et~al.\newblock {Cosmology and fundamental physics with the Euclid satellite}.
\newblock {\em Living Rev. Rel.} {\bf 2018}, {\em 21},~2.
%  \href{http://arxiv.org/abs/1606.00180}{{\normalfont
%  [arXiv:astro-ph.CO/1606.00180]}}.
[\href{http://dx.doi.org/10.1007/s41114-017-0010-3}{CrossRef}]

\bibitem[Crisostomi et~al.(2016)Crisostomi, Koyama, and
Tasinato]{Crisostomi:2016czh}
Crisostomi, M.; Koyama, K.; Tasinato, G.
\newblock {Extended Scalar-Tensor Theories of Gravity}.
\newblock {\em J. Cosmol. Astropart. Phys.} {\bf 2016}, {\em 04},~044.
%  \href{http://arxiv.org/abs/1602.03119}{{\normalfont
%  [arXiv:hep-th/1602.03119]}}.
[\href{http://dx.doi.org/10.1088/1475-7516/2016/04/044}{CrossRef}]

\bibitem[Petronikolou and Saridakis(2023)]{Petronikolou:2023cwu}
Petronikolou, M.; Saridakis, E.N.
\newblock {Alleviating the H$_{0}$ Tension in Scalar\textendash{}Tensor and
Bi-Scalar\textendash{}Tensor Theories}.
\newblock {\em Universe} {\bf 2023}, {\em 9},~397.
%  \href{http://arxiv.org/abs/2308.16044}{{\normalfont
%  [arXiv:gr-qc/2308.16044]}}.
[\href{http://dx.doi.org/10.3390/universe9090397}{CrossRef}]

\bibitem[Sakr(2023)]{Sakr:2023bms}
Sakr, Z.
\newblock {Untying the Growth Index to Relieve the \ensuremath{\sigma}$_{8}$
Discomfort}.
\newblock {\em Universe} {\bf 2023}, {\em 9},~366.
%  \href{http://arxiv.org/abs/2305.02863}{{\normalfont
%  [arXiv:astro-ph.CO/2305.02863]}}.
[\href{http://dx.doi.org/10.3390/universe9080366}{CrossRef}]

\bibitem[Sola~Peracaula et~al.(2023)Sola~Peracaula, Gomez-Valent,
de~Cruz~Perez, and Moreno-Pulido]{SolaPeracaula:2023swx}
Sola~Peracaula, J.; Gomez-Valent, A.; de~Cruz~Perez, J.; Moreno-Pulido, C.
\newblock {Running Vacuum in the Universe: Phenomenological Status in Light of
the Latest Observations, and Its Impact on the \ensuremath{\sigma}$_{8}$ and
H$_{0}$ Tensions}.
\newblock {\em Universe} {\bf 2023}, {\em 9},~262.
%  \href{http://arxiv.org/abs/2304.11157}{{\normalfont
%  [arXiv:astro-ph.CO/2304.11157]}}.
[\href{http://dx.doi.org/10.3390/universe9060262}{CrossRef}]

\bibitem[Boehmer et~al.(2023)Boehmer, Jensko, and Lazkoz]{Boehmer:2023knj}
Boehmer, C.G.; Jensko, E.; Lazkoz, R.
\newblock {Dynamical Systems Analysis of f(Q) Gravity}.
\newblock {\em Universe} {\bf 2023}, {\em 9},~166.
%  \href{http://arxiv.org/abs/2303.04463}{{\normalfont
%  [arXiv:gr-qc/2303.04463]}}.
[\href{http://dx.doi.org/10.3390/universe9040166}{CrossRef}]

\bibitem[Bouch\`e et~al.(2023)Bouch\`e, Capozziello, and
Salzano]{Bouche:2023xjw}
Bouch\`e, F.; Capozziello, S.; Salzano, V.
\newblock {Addressing Cosmological Tensions by Non-Local Gravity}.
\newblock {\em Universe} {\bf 2023}, {\em 9},~27.
%  \href{http://arxiv.org/abs/2301.01503}{{\normalfont
%  [arXiv:astro-ph.CO/2301.01503]}}.
[\href{http://dx.doi.org/10.3390/universe9010027}{CrossRef}]

\bibitem[Escamilla-Rivera and
Torres~Castillejos(2023)]{Escamilla-Rivera:2022mkc}
Escamilla-Rivera, C.; Torres~Castillejos, R.
\newblock {H0 Tension on the Light of Supermassive Black Hole Shadows Data}.
\newblock {\em Universe} {\bf 2023}, {\em 9},~14.
%  \href{http://arxiv.org/abs/2301.00490}{{\normalfont
%  [arXiv:astro-ph.CO/2301.00490]}}.
[\href{http://dx.doi.org/10.3390/universe9010014}{CrossRef}]

\bibitem[Staicova(2022)]{Staicova:2022zuh}
Staicova, D.
\newblock {DE Models with Combined H$_{0}$ \textperiodcentered{} r$_{d}$ from
BAO and CMB Dataset and Friends}.
\newblock {\em Universe} {\bf 2022}, {\em 8},~631.
%  \href{http://arxiv.org/abs/2211.08139}{{\normalfont
%  [arXiv:astro-ph.CO/2211.08139]}}.
[\href{http://dx.doi.org/10.3390/universe8120631}{CrossRef}]

\bibitem[Nesseris(2023)]{Nesseris:2022hhc}
Nesseris, S.
\newblock {The Effective Fluid approach for Modified Gravity and its
applications}.
\newblock {\em Universe} {\bf 2023}, {\em 9},~13.
%  \href{http://arxiv.org/abs/2212.12768}{{\normalfont
%  [arXiv:astro-ph.CO/2212.12768]}}.
[\href{http://dx.doi.org/10.3390/universe9010013}{CrossRef}]

\bibitem[Oikonomou et~al.(2022)Oikonomou, Tsyba, and Razina]{Oikonomou:2022tjm}
Oikonomou, V.K.; Tsyba, P.; Razina, O.
\newblock {Probing Our Universe\textquoteright{}s Past Using
Earth\textquoteright{}s Geological and Climatological History and Shadows of
Galactic Black Holes}.
\newblock {\em Universe} {\bf 2022}, {\em 8},~484.
%  \href{http://arxiv.org/abs/2209.04669}{{\normalfont
%  [arXiv:gr-qc/2209.04669]}}.
[\href{http://dx.doi.org/10.3390/universe8090484}{CrossRef}]

\end{thebibliography}

\documentclass[universe,editorial,accept,pdftex,moreauthors]{Definitions/mdpi} 

\firstpage{1} 
\pubvolume{10}
\issuenum{4}
\articlenumber{184}
\pubyear{2024}
\copyrightyear{2024}
\datereceived{3 April 2024} 
\dateaccepted{8 April 2024} 
\datepublished{18 April 2024} 
\hreflink{https://doi.org/10.3390/\linebreak universe10040184} 

\Title{Special Issue on Modified Gravity Approaches to the Tensions of $\Lambda$CDM: Goals and Highlights}

\TitleCitation{Special Issue on Modified Gravity Approaches to the Tensions of $\Lambda$CDM: Goals and Highlights}

\Author{{Eleonora Di Valentino} {$^{1}$}\orcidA{}, Leandros Perivolaropoulos {$^{2,}$}{*}\orcidB{} and Jackson Levi Said {$^{3}$}\orcidC{}}

\AuthorNames{Eleonora Di Valentino, Leandros Perivolaropoulos and Jackson Levi Said}

\AuthorCitation{{Di Valentino}, E.; Perivolaropoulos, L.; Said, J.L.}

\address{%
$^{1}$ \quad School of Mathematics and Statistics, University of Sheffield, {Sheffield S10 2AH}, UK; 
 e.divalentino@sheffield.ac.uk\\
$^{2}$ \quad Department of Physics, University of Ioannina, {45110 Ioannina}, Greece\\
$^{3}$ \quad Institute of Space Sciences and Astronomy, University of Malta, {2080 Msida}, 
Malta; jackson.said@um.edu.mt}

\corres{Correspondence: leandros@uoi.gr}

\abstract{The Special Issue on \href{https://www.mdpi.com/journal/universe/special_issues/Gravit}{"Modified Gravity Approaches to the Tensions of $\Lambda$CDM"} in the Universe journal tackles significant challenges faced by the $\Lambda$CDM model, including discrepancies in the Hubble constant, growth rate of structures, and cosmological anisotropies. These issues suggest foundational cracks in the model, raising questions about the validity of General Relativity, dark energy, and cosmological principles at large scales. This collection brings together leading researchers to delve into Modified Gravity theories as potential solutions. Covering approaches from Scalar-Tensor theories to $f(R,T)$ gravity and beyond, each contribution presents innovative research aimed at addressing the limitations of the $\Lambda$CDM model. This Special Issue not only highlights the theoretical and empirical strengths of Modified Gravity models but also opens avenues for future investigations, emphasizing the synergy between theoretical advancements and observational evidence to deepen our cosmological understanding.}

\keyword{cosmology; $\Lambda$CDM model; modified gravity; Hubble tension; dark energy}

\begin{document}

\setcounter{section}{0} 
\section{Introduction}
\label{sec:introduction}

The standard cosmological model, known as $\Lambda$CDM, has been remarkably successful in providing a coherent and predictive framework for understanding the Universe's evolution, its large-scale structure, and cosmic microwave background (CMB) radiation~\cite{Planck:2018vyg,ACT:2023dou,SPT-3G:2022hvq,eBOSS:2020yzd,KiDS:2020suj,DES:2021wwk}. Central to this model are the cosmological constant $\Lambda$, representing dark energy responsible for the accelerated expansion of the Universe, and cold dark matter (CDM), which accounts for the gravitational scaffolding underlying galaxy formation and evolution. Despite its triumphs, the $\Lambda$CDM model is not without its challenges~\cite{Perivolaropoulos:2021jda,Abdalla:2022yfr}. In recent years, high-precision cosmological observations have revealed a series of tensions that question the completeness of the $\Lambda$CDM paradigm. These tensions---most notably, the discrepancy in the Hubble constant ($H_0$) measurements from local and early Universe observations~\cite{Verde:2019ivm,DiValentino:2020zio,DiValentino:2021izs,Vagnozzi:2023nrq,Dainotti:2023yrk,Akarsu:2024qiq,Kamionkowski:2022pkx,Dainotti:2021pqg,Perivolaropoulos:2021jda,Jedamzik:2020zmd,Vagnozzi:2021tjv}, the growth rate of cosmic structures~\cite{DiValentino:2020vvd,Philcox:2021kcw,Heisenberg:2022gqk,Benisty:2020kdt,Kilo-DegreeSurvey:2023gfr,KiDS:2020ghu,Joudaki:2019pmv,Troster:2019ean}, and the scale of cosmological \mbox{anisotropies~\cite{Aluri:2022hzs,Secrest:2020has,Uzan:2008qp,McConville:2023xav,Watkins:2023rll,Cowell:2022ehf,Dam:2022wwh,Krishnan:2022qbv,Bengaly:2024ree,Ebrahimian:2023svi,Perivolaropoulos:2023tdt}}---suggest potential shortcomings in our understanding of the fundamental components and laws governing the Universe.

The emergence of these tensions has spurred a vibrant discourse within the cosmological community, giving rise to the motivation behind this Special Issue on ``Modified Gravity Approaches to the Tensions of $\Lambda$CDM''. These inconsistencies may herald the advent of a new standard cosmological model, one rooted in physics beyond the current paradigm. As we stand on the brink of potentially revolutionary discoveries, it becomes imperative to explore avenues beyond the conventional framework, questioning the validity of general relativity (GR) on cosmological scales and the nature of dark energy and dark matter.

Modified gravity theories offer a well-motivated and generic theoretical framework for extending and potentially supplanting the standard $\Lambda$CDM model. These theories, which range from scalar--tensor theories to more exotic formulations such as $f(R,T)$ gravity and non-metricity gravity, propose alterations or extensions to GR that can naturally account for the accelerated expansion of the Universe without resorting to the cosmological constant or {elucidating the dynamics of cosmic structure formation in novel ways}. 
Their appeal lies not only in their ability to address the existing tensions~\cite{Bahamonde:2021gfp,Adi:2020qqf,Marra:2021fvf,Odintsov:2020qzd,Ballardini:2020iws,Wang:2020zfv,Skara:2019usd,Escamilla-Rivera:2019ulu,Cai:2019bdh,Kazantzidis:2019nuh,Kazantzidis:2018rnb,Nunes:2018xbm,Gangopadhyay:2022bsh,Zumalacarregui:2020cjh,Benevento:2022cql,Braglia:2020auw,Farhang:2020sij,Heisenberg:2020xak,Ballesteros:2020sik,FrancoAbellan:2023gec} but also in their potential to enrich our theoretical landscape with new physics that could resolve longstanding puzzles in cosmology~\cite{CANTATA:2021ktz,Langlois:2017dyl,Nojiri:2017ncd,Joyce:2016vqv,Joudaki:2020shz,Koyama:2015vza,Kobayashi:2019hrl,Ishak:2018his,Sakstein:2017xjx,Amendola:2016saw,Crisostomi:2016czh}.

{This} \href{https://www.mdpi.com/journal/universe/special_issues/Gravit}{Special Issue} aims to delve into the heart of these debates, presenting a collection of cutting-edge research that explores the theoretical viability, empirical implications, and observational constraints of modified gravity theories. Through this collective endeavor, we seek to illuminate the pathways toward a deeper understanding of the cosmos, guided by the principle that the resolution of the $\Lambda$CDM tensions could unveil new facets of our Universe and lay the groundwork for a new standard model of cosmology.


\section{Overview of the Published Articles}
\label{sec:overview}

Maria Petronikolou and Emmanuel N. Saridakis's article (contribution 1~\cite{Petronikolou:2023cwu}) delves into scalar--tensor- and bi-scalar--tensor-modified theories of gravity as potential frameworks for alleviating the Hubble tension. By selecting models with a unique shift-symmetric friction term, their work demonstrates how these theories can significantly mitigate the discrepancy between local and early Universe measurements of the Hubble constant, offering new pathways for resolving this longstanding cosmological puzzle.

Ziad Sakr's research (contribution 2~\cite{Sakr:2023bms}) focuses on untangling the $\sigma_8$ discomfort by independently analyzing the matter fluctuation parameter $\sigma_8$ and the growth index $\gamma$. His innovative approach treats $\sigma_8$ as a free parameter, distinct from its traditional derivation, revealing how this separation can lead to more accurate constraints on cosmological parameters and offering a novel perspective on the growth tension within the $\Lambda$CDM~framework.

Joan Solà Peracaula and colleagues (contribution 3~\cite{SolaPeracaula:2023swx}) investigate the Running Vacuum Model (RVM) as a dynamical alternative to the cosmological constant. Their extensive analysis, supported by the latest observational data, showcases the RVM's ability to provide a compelling fit to cosmic phenomena while potentially resolving both the $\sigma_8$ and $H_0$ tensions, marking a significant step towards a dynamic understanding of dark~energy.

Christian Böhmer, Erik Jensko, and Ruth Lazkoz (contribution 4~\cite{Boehmer:2023knj}) employ dynamical systems analysis to explore $f(Q)$ gravity's implications for cosmological evolution. Their study elucidates how modifications in the gravity sector can lead to viable cosmological models that offer new insights into the accelerated expansion of the Universe, challenging the conventional $\Lambda$CDM model with a fresh theoretical perspective.

Mayukh R. Gangopadhyay and colleagues (contribution 5~\cite{Gangopadhyay:2022bsh}) present a scenario of large-scale modification of gravity without invoking extra degrees of freedom. Their model, incorporating interactions between baryonic and dark matter, offers a unified framework to address both the late-time acceleration of the Universe and the Hubble tension, suggesting a seamless integration of dark sector phenomena within modified gravity theories.

Filippo Bouché, Salvatore Capozziello, and Vincenzo Salzano (contribution 6~\cite{Bouche:2023xjw}) tackle cosmological tensions through the lens of non-local gravity. By revisiting the foundations of gravitational interaction, their work highlights how non-local modifications can reconcile discrepancies in $\Lambda$CDM predictions, opening the door to alternative models of dark energy and gravity.

\textls[-5]{Celia Escamilla-Rivera and Rubén Torres Castillejos (contribution 7~\cite{Escamilla-Rivera:2022mkc}) explore the Hubble tension using supermassive black hole shadows data. Their innovative approach leverages high-resolution astrophysical observations to infer $H_0$, demonstrating the potential of black hole shadows as novel probes in cosmology and offering a fresh avenue for resolving observational tensions.}

Denitsa Staicova's contribution (contribution 8~\cite{Staicova:2022zuh}) examines dynamical dark energy models in light of combined $H_0 \cdot r_d$ measurements. By analyzing the multiplication of the Hubble parameter and sound horizon scale as a single parameter, Staicova provides new insights into dark energy dynamics, suggesting that dynamical models could offer a resolution to the Hubble tension.

Savvas Nesseris (contribution 9~\cite{Nesseris:2022hhc}) reviews the effective fluid approach as a versatile framework for representing a wide array of modified gravity models. By treating modified gravity effects as contributions from an effective dark energy fluid, Nesseris provides a unified analysis tool for comparing theoretical predictions with observational data. This approach not only facilitates the examination of modified gravity's role in cosmic acceleration but also provides a systematic method for confronting these models with the growth rate of structure and the expansion history of the Universe, offering a bridge between theory and observation.

V. K. Oikonomou, Pyotr Tsyba, and Olga Razina (contribution 10~\cite{Oikonomou:2022tjm}) offer a novel perspective by exploring how Earth's geological and climatological history, alongside the shadows of galactic black holes, might reveal insights into our Universe's evolution. Their interdisciplinary approach highlights potential signatures of past pressure singularities and their implications for the early Universe's dynamics. This intriguing exploration underscores the untapped potential of combining astrophysical, geological, and climatological data to inform and constrain cosmological models, opening up new avenues for understanding the Universe's past and the nature of gravity.

Sunny Vagnozzi (contribution 11~\cite{Vagnozzi:2023nrq}) presents a compelling argument that early-time physics modifications alone may be insufficient to resolve the Hubble tension. By synthesizing evidence across seven hints derived from cosmological observations and theoretical considerations, Vagnozzi's work critically assesses the landscape of proposed solutions to the Hubble tension. This contribution emphasizes the necessity for a comprehensive approach that accounts for both early- and late-time Universe physics, challenging the cosmological community to broaden the scope of theoretical explorations in search of a more complete resolution to one of modern cosmology's most pressing puzzles.

Together, these eleven contributions exemplify this Special Issue's commitment to exploring the frontiers of modified gravity theories as viable alternatives or extensions to the $\Lambda$CDM model. Each article advances the dialogue within the cosmological community, offering fresh insights, innovative methodologies, and compelling arguments that collectively enrich our understanding of the Universe's fundamental nature and its governing~laws.

\section{Conclusions}
\label{sec:conclusion}

The Special Issue ``Modified Gravity Approaches to the Tensions of $\Lambda$CDM'' represents a significant stride toward addressing some of the most perplexing challenges in modern cosmology. Through a diverse collection of contributions, this endeavor has not only contributed to the illumination of the intricacies of the tensions within the $\Lambda$CDM model, but has also underscored the potential of modified gravity theories to pave the way for groundbreaking discoveries. Each article, in its unique capacity, has contributed to a deeper understanding of the Universe's fundamental laws, showcasing the richness and vitality of theoretical innovation in the face of empirical puzzles.

The collective progress made through this Special Issue is a testament to the importance of continued exploration and open-mindedness in the scientific quest to understand the cosmos. The detailed examinations of modified gravity theories presented here underscore the necessity of extending our theoretical frameworks beyond the confines of general relativity and the standard cosmological model. These theories offer not just solutions to specific observational tensions, but also a broader perspective on gravity and its role in the evolution of the Universe.

Moreover, this Special Issue stands as a potential milestone in the discovery of new physics. The tensions within the $\Lambda$CDM model may indeed signal the limitations of our current understanding, pointing toward an underlying reality that is more complex and nuanced than previously thought. In this context, modified gravity theories emerge not merely as alternatives, but as harbingers of a new standard model of cosmology. They represent the most fundamental approach to addressing the cosmological tensions, weaving together the empirical anomalies with theoretical insights to sketch a more accurate and comprehensive picture of the Universe.

The journey toward resolving the enduring tensions within the $\Lambda$CDM model and unveiling the new physics that will undoubtedly reshape our cosmological paradigm is far from over. However, the contributions of this Special Issue mark critical waypoints on this journey. They invite the scientific community to reconsider the foundations of cosmology, to embrace the uncertainties and anomalies, not as mere nuisances, but as beacons guiding us toward a more profound understanding of the Universe.

As we continue to probe the depths of the cosmos with ever more sophisticated tools and technologies, the insights gleaned from modified gravity theories will undoubtedly play a special role. They offer a promising path forward, one that harmonizes the elegance of theoretical physics with the empirical realities of the Universe we strive to understand. In this endeavor, the Special Issue ``Modified Gravity Approaches to the Tensions of $\Lambda$CDM'' stands as both a milestone and a beacon, illuminating the path towards the next standard model of cosmology and the new physics that will underpin it.
\vspace{6pt}

\conflictsofinterest{The authors declare no conflicts of interest.}


\begin{adjustwidth}{-\extralength}{0cm}
\reftitle{References}

\PublishersNote{}
\end{adjustwidth}


\end{document}